OpenStudy (anonymous):

Simplify (x^2/3)^4/5

OpenStudy (jdoe0001):

\(\Large \bf { (a^n)^m\implies a^{n\cdot m} \\ \quad \\ \quad \\ \left(x^{\frac{2}{3}}\right)^{\frac{4}{5}}\implies x^{\frac{2}{3}\cdot \frac{4}{5}} } \)